OpenStudy (anonymous):

find the zeros of the polynomial and state the multiplicity: f(x)=x^3-x^2-4x+4

myininaya (myininaya):

How do you find the possible rational zeros?

OpenStudy (anonymous):

that's what I don't understand :/

myininaya (myininaya):

Can you answer these questions? Find the factors of the constant term. Find the factors of the coefficient of the term with the highest exponent.

myininaya (myininaya):

Find the factors of the constant term means find the factors of 4 since 4 is the constant term. So what are the factors of 4?

OpenStudy (anonymous):

1, 2, and 4?

myininaya (myininaya):

well plus or minus but okay. Now what are the factors of the coefficient of the term with the highest exponent. This means I'm looking at 1 since 1 is in front of x^3. What are the factors of 1?

OpenStudy (anonymous):

1

myininaya (myininaya):

well plus or minus but yeah

OpenStudy (anonymous):

oh okay so +/- 1

myininaya (myininaya):

We define the possible rational zeros as the set containing the factors of the constant term/ the factors of the coefficient of the term with highest exponent. So this means we have \[\text{ possible rational zeros } = \{ {\frac{\pm 1}{\pm 1} , \frac{\pm 2}{\pm 1}, \frac{\pm 4}{\pm 1} }\}\]

myininaya (myininaya):

or just simply \[\{ \pm 1 , \pm 2 , \pm 4\} \]

myininaya (myininaya):

Now possible doesn't mean actual.

myininaya (myininaya):

This just narrows are search.

myininaya (myininaya):

So we have f(x)=x^3-x^2-4x+4 . We try some of our possible rational zeros to see which gives us 0.

myininaya (myininaya):

What value in our possible rational set do you want to try and see if it works?

OpenStudy (anonymous):

would we do 1 since it is first?

myininaya (myininaya):

You can but I was asking you where you want to start.

myininaya (myininaya):

So if we plug in 1 do we get 0?

OpenStudy (anonymous):

am I plugging 1 in for x?

myininaya (myininaya):

We are trying to find for what values of x is f zero so yes.

OpenStudy (anonymous):

okay by plugging 1 in for x it does equal 0

myininaya (myininaya):

\[f(1)=1^3-1^2-4(1)+4 \] is this zero? You are right the answer is yes. x=1 is a zero since f will be 0 when x=1 If x=1 is a zero then x-1 is a factor of f. Now divide f by x-1. You may use synthetic or long polynomial division.

OpenStudy (anonymous):

what would my equation be for that? do I just plug in x-1?

myininaya (myininaya):

for f?

myininaya (myininaya):

like you are dividing f by x-1 to find the other factors of f.

myininaya (myininaya):

remember f is x^3-x^2-4x+4

myininaya (myininaya):

f/(x-1)=?

myininaya (myininaya):

\[\frac{x^3-x^2-4x+4}{x-1} =?\]

OpenStudy (anonymous):

\[0+ \frac{ 4x }{ x-1 }\]

myininaya (myininaya):

No that is way off. I'm sorry.

myininaya (myininaya):

Do you know how to do long division with polynomials?

OpenStudy (anonymous):

I just don't understand how to do this I suppose

myininaya (myininaya):

Have you discussed long division? I could show you an easier way but doesn't always work out so well like polynomial division does.

OpenStudy (anonymous):

I tried to do the problem in synthetic division but it just said it's not possible

myininaya (myininaya):

it is possible

myininaya (myininaya):